The length of a rectangle is 5 centimeters less than twice its width. The perimeter of the rectangle is 26 cm. What are the dimensions of the rectangle?
Answer by checkley77(12844)

L=2W-5
2L+2W=26
2(2W-5)+2W=26
4W-10+2W=26
6W=26+10
6W=36
W=36/6
W=6 ans. for the width.
L=2*6-5=12-5=7 ANS. FOR THE LENGTH.
PROOF:
2*7+2*6=26
14+12=26
26=26